# Neural Beam Field for Spatial Beam RSRP Prediction

### v1

@@ IV-A Validation of Analytical Formulas
To assess the validity of the derived theoretical model under realistic channel conditions, we simulate an outdoor scenario using the QuaDRiGa toolbox. For each beam and receiver location, we compute the theoretical mean and standard deviation of the RSRP under given MCPPs. In the MC simulations, we average over $N_{\mathrm{mc}}=200$ independent channel realizations, apply the DFT beamforming weights in each realization, and compute the empirical means and standard deviation of the instantaneous received power.
To validate the accuracy of our theoretical model, we first compare the mean and standard deviation of the beam RSRP predicted by our analytical formulas with those obtained from the MC simulations. For a given location, as shown in Fig. 3, the results show good agreement across different beam configurations. Furthermore, we compare the spatial distribution of the mean RSRP over the entire map, as depicted in Fig. 4, which confirms the consistency between the theoretical predictions and simulation results.
@@
Figure 4: Comparions of spatial mean RSRP heatmap for a representative beam over the entire map region, obtained by either MC or formulas.
@@ IV-B Effectiveness of NBF
For NBF training, we collect ray-traced user channels and RSRP samples on a site-specific map region. Each sample consists of the UE position, antenna panel and beam configurations, together with the associated performance label. We use an $80/20$ split of dataset for training and validation, and optimize the Smooth-$L_{1}$ loss on the mean RSRP using the Adam optimizer and OneCycleLR learning rate schedule.
Fig. 5 compares the training and validation loss curves for three approaches, i.e., 1) NBF with end-to-end learning from RSRP data, 2) NBF fine-tuning after MCPP-prior aided pretraining, and 3) an MLP baseline, also with end-to-end learning from RSRP data. First, it can be seen that NBF outperforms the MLP baseline in terms of convergence speed and prediction accuracy, both under end-to-end training. Second, the MCPP-prior aided pretraining helps tame the NBF into better initial points and further reduce both the training and validation losses, thus improving prediction accuracy and generalization.
Next, we evaluate the effectiveness of NBF through an illustrative example in Fig. 6, which presents spatial mean RSRP heatmaps for a representative beam under ground truth, pretrained NBF, and MLP. It can be seen that the NBF prediction closely follows the ground-truth distribution and captures fine spatial details, while the MLP output appears overly smooth and misses local details.
Figure 5: Training and validation loss curves for end-to-end NBF training (left), NBF fine-tuned after MCPP prior pretraining (middle), and an MLP baseline trained end-to-end (right).
Figure 6: Spatial mean RSRP heatmap comparison for a representative beam: ground truth (left), NBF prediction (center), and MLP prediction (right).
In terms of quantitative metrics, we evaluate the MAE between the predicted and ground-truth mean RSRPs (normalized to transmit power $P_{\text{t}}$) in dB scale, as well as the storage requirement in mega-bytes (MB) for CKM tables or DNN parameters, under three different schemes. Our NBF architecture under end-to-end training achieves MAE of $3.547$ dB while requiring only $8.21$ MB of storage. The MLP baseline recordes MAE of $4.105$ dB and requires $10.57$ MB of storage. These results show that NBF achieves higher prediction accuracy with lower storage requirement. When prior information about MCPP is available, we apply the MCPP prior to both the NBF pretraining and the IDW baseline, both of which utilize our whitebox analytical formulas to map MCPP into beam RSRP statistics. In this case, the MCPP-prior aided pretraining helps reduce its MAE from $3.547$ dB to $2.416$ dB, while IDW yields MAE of $3.328$ dB and consumes $27.86$ MB of storage. As a result, NBF outperforms IDW in fitting quality and storage efficiency.

### v2

@@ IV-A Validation of Analytical Formulas
We simulate an outdoor scenario using the QuaDRiGa toolbox, which implements 3GPP 38.901 channel models. In the MC simulations, we average over $N_{\mathrm{mc}}=200$ independent channel realizations, apply the DFT beamforming weights in each realization, and compute the empirical mean and standard deviation of the instantaneous received power. For a given UE location, the results are plotted in Fig. 3, with good match with our analytical formulas across different beam configurations. Furthermore, good consistency has also been observed across the entire map, which is omitted due to space limit.
@@ IV-B Effectiveness of NBF
For NBF training, we collect ray-traced user channels and RSRP samples on a site-specific map. Each sample consists of the UE position, antenna panel and beam configurations, together with the associated performance label. We use an $80/20$ split of dataset for training and validation, and optimize the Smooth-$L_{1}$ loss on the mean RSRP using the Adam optimizer and OneCycleLR learning rate schedule.
Firstly, We evaluate the effectiveness of NBF through an illustrative example in Fig. 4, which presents spatial mean RSRP heatmaps for a representative beam under ground truth, pretrained NBF, and MLP. It can be seen that the NBF prediction closely follows the ground-truth distribution and captures fine spatial details, while the MLP output appears overly smooth and misses local details.
Figure 4: Spatial mean RSRP heatmap comparison for a given beam: ground truth (left), NBF (center), and MLP (right).
Furthermore, quantitative results are compared in Table I, including mean absolute error (MAE) of mean RSRP prediction, model storage requirement in Mega-Bytes (MB), and computation/inference time in milliseconds (ms). It can be seen that NBF provides the best overall performance. Although the RSRP-based IDW method is faster in inference, its prediction accuracy remains inferior to our NBF model even with larger storage (and correspondingly more RSRP measurements).
TABLE I: Performance Comparison of NBF and Baseline Models
Model MAE (dB) Storage (MB) Time (ms)
NBF (Proposed) 1.218 16.4 0.273
MLP (Baseline) 3.423 16.5 0.212
IDW (MCPP-based) 2.996 27.0 3.365
IDW (RSRP-based) 1.937 55.3 0.062
Finally, we evaluate the MAE of mean RSRP prediction under different number of available samples, as shown in Fig. 5, under ray-traced deterministic channel (left) and hybrid channel with add-on random MCPP components (right). In the deterministic scenario, both IDW methods perform well (better than MLP), with their accuracy improving as the number of samples grows. However, their performance degradates significantly in hybrid channels due to random channel factors, especially for the MCPP-based IDW method which relies on accurate MCPP priors. In comparison, our NBF methods provides the overall best performance under both settings. In particular, the MCPP-prior guided NBF performs consistently the best even under random perturbation of MCPP, which suggests that our PaC strategy effectively learns the physical groundings underneath the RSRP measurements and thus provides accurate/robust beam RSRP predictions under diverse practical settings.
Figure 5: MAE comparison under ray-traced deterministic channel (left) and hybrid channel with add-on random MCPP components (right).